OpenStudy (anonymous):

Find the slope and the y-intercept of the line. 7x=9y+81

OpenStudy (anonymous):

re-write it as: 9y=7x-81 y=7/9 x -9 the slope is the coefficient in a front of x.. what is slope equal to?

OpenStudy (anonymous):

y-intercept => x=0 if x=0 what is y=?

OpenStudy (anonymous):

the slope is 7/9 the y intercept is (-9,0). you need to change it to y=mx+b (slope-intercept form) m is the slope and b is your y intercept.

OpenStudy (anonymous):

slope m=7/9 y-intercept: y=-9 or like Jerni wrote: (0,-9) got to go , guys...good luck!

OpenStudy (anonymous):

Thanks